OpenStudy (anonymous):

6. You spin a spinner that has eight equal-sized sections numbered one to eight. Find P(perfect square or odd).

OpenStudy (lalaly):

perfect squares from 1 to 8 are 1,4 The odd numbers are 1,3,7,5 P(perfect square or odd) = P(1, 3, 4, 5 or 7) = P(1) + P(3) + P(4) + P(5) + P(7) = 1/8 + 1/8 + 1/8 + 1/8 + 1/8 = 5/8

OpenStudy (unklerhaukus):

perfect square or odd =1-P(not square nor odd) =1-(P(2)+P(6+P(7) =1-(1/8+1/8+1/8) =1-3/8 =5/8

OpenStudy (lalaly):

@UnkleRhaukus 1-(p(2)+p(6)+p(8))

OpenStudy (anonymous):

Favourable cases ={1,3,4,5,7} P=5/8

OpenStudy (unklerhaukus):

yeah @lalaly, that is what i ment on line 2